# Quantum Kernels for Symbolic Sequences Classification

### v2

@@ 2.3 Classification Problem for Stochastic Languages
We consider symbolic time series classification problem for a stochastic process language $L$ and a finite set of class labels $C=\bigl{\{}c_{1},\cdots,c_{k}\bigr{\}}$. It is assumed that there exists an unknown probabilistic classification map $\mathcal{P}:L\rightarrow P[C]$ assigning to each sequence $\mathbf{y}\in L$ a distribution over the class labels. We consider two types of probabilistic classification maps. The first classification map depends solely on the structure of a given sequence $\mathbf{y}$ and is defined by a mapping function $p_{s}$ estimating the probability of each class label $c$ for $\mathbf{y}$:
$p_{s}(\mathbf{y},c)=P[c|\mathbf{y}],\mathbf{y}\in L,c\in C.$ (11)
We refer to classification problems with mapping function (11) as structural. For such problems, the class of a sequence is directly associated with its structure. Examples of structural problems include the classification of protein sequences into functional families [41], detection of fraudulent financial transactions [42], and monitoring for threats in the cybersecurity domain [43] ], to name a few.
In another scenario the classification map depends on the probabilities and structure of the future evolution or continuations with length $k$ of a sequence $\mathbf{y}$ and is defined as:
$p_{p}(\mathbf{y},k,c)=\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}]P[c|\mathbf{z}]),\mathbf{y}\in L,c\in C,$ (12)
where $P[\mathbf{z}|\mathbf{y}]$ is the probability to observe $\mathbf{z}$ as continuation of $\mathbf{y}$. These types of classification problems are known as predictive. Predictive classification problems find applications in various domains such as [44], stock market prediction [45], health monitoring [46], energy load forecasting [47], and traffic flow prediction [48].
A training data sample is defined as follows:
$S=\bigl{\{}(\mathbf{y}_{i},c_{i}):\mathbf{y}_{i}\in Y,c_{i}\in C,i\in\left[1,l\right]\bigr{\}},$
where $Y=\bigl{\{}\mathbf{y}:\mathbf{y}\sim D^{L}_{i},i\in\left[1,l\right]\bigr{\}}$ is a set of finite length sequences sampled from the distributions $D^{L}_{i}$ of $L$.
The goal of the classification problem is to learn an approximation $\mathcal{\hat{P}}$ of the unknown probabilistic classification map.
The performance of the approximated mapping function $\hat{p}(\mathbf{y},c)$ is typically evaluated using probabilistic metrics such as log-likelihood, Brier score, or other probabilistic measures. These metrics assess the model’s ability to predict the probability distribution over class labels and quantify the quality of the stochastic classification results.
@@ 3 Quantum Hidden Markov Models
If the operation $\mathcal{T}$ is applied $t$ times starting at the initial state a time series $\mathbf{y}=y_{1}\ldots y_{t}$ will be observed with probability
@@ 4.1 Predictive Generative Kernels
The predictive similarity measure is defined by the stochastic distance between the expected future evolutions of the sequences. Since the evolution of a Markovian process depends only on its current state, it is easy to verify that if two sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ define the same quantum sate $\rho$ of the model (23):
$\rho=\frac{T_{\mathbf{y^{1}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{1}}|\rho_{0}\bigr{]}}}=\frac{T_{\mathbf{y^{2}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{2}}|\rho_{0}\bigr{]}}},$
then the sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ are considered equivalent in the sense, that they define the same future distributions of the observables.
@@
For every sequences $\mathbf{y}^{1}$ and $\mathbf{y}^{2}$ in the example domain $L^{\mathbf{Q}}$ of a QHMM $\mathbf{Q}$ a predictive kernel $\kappa_{p}:L^{\mathbf{Q}}\times L^{\mathbf{Q}}\rightarrow\mathbb{R}$ is the function:
$\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})=\mathcal{D}\bigl{(}\phi(\mathbf{y}^{1}),\phi(\mathbf{y}^{2})\bigr{)},\mathbf{y}^{1},\mathbf{y}^{2}\in L^{Q}$ (37)
where the mapping $\phi(\mathbf{y})$ of a sequence $\mathbf{y}$ to quantum state $\rho_{\mathbf{y}}$ is defined as (23) :
$\phi(\mathbf{y})=\frac{T_{\mathbf{y}}\rho_{0}}{P\bigl{[}{\mathbf{y}|\rho_{0}\bigr{]}}},\mathbf{y}\in L^{Q},$
and $\mathcal{D}(\rho_{\mathbf{y}},\rho_{\mathbf{z}})$ is the trace distance of the quantum states (29). Since the trace distance is a symmetric, positive semi-definite function the equation (37) defines a valid kernel.
Critical feature of the kernel is that it reflects the similarity of sequences with respect of their probabilistic classification maps (12) as it is demonstrated by the following Proposition.
Proposition 2.
For any two sequences $\mathbf{y}^{1}$ and $\mathbf{y}^{2}$ the variation distance of the classification maps (12) $p_{p}(\mathbf{y}^{1},k,c)$ and $p_{p}(\mathbf{y}^{2},k,c),k>0$ is bounded by the predictive kernel distance $\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})$.
@@
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}]P[c|\mathbf{z}])-\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{2}]P[c|\mathbf{z}])\bigr{|}$ (39)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}P[c|\mathbf{z}](P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}])\bigr{|}$ (40)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=c^{*}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}])\bigr{|}$ (41)
@@
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq 2c^{*}|\Sigma^{k}|\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2}),$ (44)
where $c^{*}\leq 1$ is a constant dependent on the particular probabilistic classification map $P[c|\mathbf{z}]$.
Another predictive quantum kernel can be defined by the Bures metric between two quantum states defined as
@@
Since the Bures metric is symmetric and positive semi-definite the following equality defines a valid kernel:
$\kappa_{p}(\mathbf{y},\mathbf{z})=\mathcal{B}\bigl{(}\phi(\mathbf{y}),\phi(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}.$ (45)
@@ 4.2 Structural Generative Kernels
$\kappa_{s}(\mathbf{y},\mathbf{z})=\mathcal{D}\bigl{(}\phi_{s}(\mathbf{y}),\phi_{s}(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}$ (47)
@@
$\kappa_{s}(\mathbf{y},\mathbf{z})=\mathcal{B}\bigl{(}\phi_{s}(\mathbf{y}),\phi_{s}(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}$ (48)
@@ 5.3 Quantum Kernels vs Classical Kernels
where $\left\lVert(\mathbf{y}^{1}-\mathbf{y}^{2})\right\rVert$ is the Euclidean distance between the sequences, and $\sigma$ is a parameter. In the quantum implementation of the RBF kernel we replaced the the Euclidean distance with the quantum predictive kernel $\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})$ (37),(45):
$K_{q}(\mathbf{y}^{1},\mathbf{y}^{2})=\operatorname{exp}\left(-\frac{\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})}{2\sigma^{2}}\right),$
and the quantum structural kernel $\kappa_{s}(\mathbf{y}^{1},\mathbf{y}^{2})$ (47), (48):
$K_{q}(\mathbf{y}^{1},\mathbf{y}^{2})=\operatorname{exp}\left(-\frac{\kappa_{s}(\mathbf{y}^{1},\mathbf{y}^{2})}{2\sigma^{2}}\right),$
@@ 6 Quantum Kernels in Quantum Circuits Computing Model
In the case of more general feature maps producing mixed states we define a kernel function $\kappa$ as follows:
$\kappa(\mathbf{a},\mathbf{b})=1-\mathcal{D}(\rho_{\mathbf{a}},\rho_{\mathbf{b}}),$
where $\mathcal{D}$ is a distance between two general quantum states, represented by their density matrices.

### v3

# Quantum Time Series Similarity Measures and Quantum Temporal Kernels

@@ 2.3 Classification Problem for Stochastic Languages
We consider the symbolic time series classification problem for a stochastic language $L$ and a finite set of class labels $C=\{c_{1},\ldots,c_{k}\}$. It is assumed that there exists an unknown probabilistic classification map $\mathcal{P}:L\rightarrow P[C]$, which assigns to each sequence $\mathbf{y}\in L$ a probability distribution over the class labels.
We distinguish between two types of probabilistic classification maps:
Structural classification maps.
These maps, which we call structural, assign class probabilities based solely on the structure of the observed sequence:
$p_{s}(\mathbf{y},c)=P[c\mid\mathbf{y}],\quad\mathbf{y}\in L,\;c\in C.$ (11)
In structural classification, the class of a sequence depends only on its observed pattern. Examples include protein family classification based on amino acid sequences [41], detection of anomalous financial transactions [42], and intrusion detection in cybersecurity [43].
Predictive classification maps.
The second type of classification maps assign class probabilities depending on the expected future evolution of the observed sequence. For a fixed horizon $k>0$, the predictive classification map is defined as:
$p_{p}(\mathbf{y},c)=\sum_{\mathbf{z}\in\Sigma^{k}}P[\mathbf{z}\mid\mathbf{y}]\cdot P[c\mid\mathbf{yz}],\quad\mathbf{yz}\in L,\;c\in C,$ (12)
where $P[\mathbf{z}\mid\mathbf{y}]$ denotes the probability of observing continuation $\mathbf{z}$ of length $k$ after observing $\mathbf{y}$, and $P[c\mid\mathbf{yz}]$ is the class probability conditioned on $\mathbf{yz}$.
We refer to classification problems defined by (12) as predictive. These problems arise in domains where future dynamics influence class membership, such as weather forecasting [44], stock market trend prediction [45], patient health monitoring [46], energy load forecasting [47], and traffic flow analysis [48].
Training and learning.
A training dataset consists of labeled sequence samples:
$S=\left\{(\mathbf{y}_{i},c_{i})\mid\mathbf{y}_{i}\in Y,\;c_{i}\in C,\;i\in[1,\ell]\right\},$
where $Y=\left\{\mathbf{y}_{i}\sim D^{L}_{i}\right\}$ is a set of finite sequences sampled from respective distributions $D^{L}_{i}$ over $L$. The goal of the classification problem is to learn an approximation $\widehat{\mathcal{P}}$ to the true classification map $\mathcal{P}$.
Evaluation.
The quality of the learned classifier $\hat{p}(\mathbf{y},c)$ is typically evaluated using probabilistic scoring rules such as the Brier score, negative log-likelihood, or proper scoring metrics that assess the predictive distribution over labels. These measures quantify the fidelity of the learned stochastic classification behavior.
@@ 3 Quantum Hidden Markov Models
If the operation $\mathcal{T}$ is applied $t$ times starting at the initial state a sequence $\mathbf{y}=y_{1}\ldots y_{t}$ will be observed with probability
@@ 4.1 Predictive Generative Kernels
The predictive similarity measure is defined by the stochastic distance between the expected future evolutions of the sequences. Since the evolution of a Markovian process depends only on its current state, it is easy to verify that if the quantum states $\rho_{1}$ and $\rho_{2}$ defined by any two sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ (23):
$\rho_{1}=\frac{T_{\mathbf{y^{1}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{1}}|\rho_{0}\bigr{]}}}$
$\rho_{2}=\frac{T_{\mathbf{y^{2}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{2}}|\rho_{0}\bigr{]}}},$
are the same, then the sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ are considered equivalent in the sense, that they define the same future distributions of the observables.
@@
Let $\mathbf{Q}$ be a Quantum Hidden Markov Model (QHMM) with domain of observable sequences $L^{\mathbf{Q}}$. For any two sequences $\mathbf{y}^{1},\mathbf{y}^{2}\in L^{\mathbf{Q}}$, the predictive generative quantum kernel $\kappa_{p}:L^{\mathbf{Q}}\times L^{\mathbf{Q}}\to\mathbb{R}$ is defined as:
$\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})=\exp\left(-\mathcal{D}\left(\phi(\mathbf{y}^{1}),\phi(\mathbf{y}^{2})\right)\right),$ (37)
where $\phi:L^{\mathbf{Q}}\to\mathcal{D}(\mathcal{H})$ maps each sequence $\mathbf{y}$ to a normalized quantum state $\rho_{\mathbf{y}}$(23):
$\phi(\mathbf{y})=\frac{T_{\mathbf{y}}\rho_{0}}{P[\mathbf{y}\mid\rho_{0}]},$ (38)
with $T_{\mathbf{y}}$ denoting the sequence of Kraus operators associated with $\mathbf{y}$, $\rho_{0}$ the initial state of the system, $P[\mathbf{y}\mid\rho_{0}]$ the probability of observing sequence $\mathbf{y}$ given the initial state, and $\mathcal{D}(\rho_{\mathbf{y}},\rho_{\mathbf{z}})$ is the trace distance between the quantum states (29).
Since the trace distance is a conditionally negative definite (CND) metric on quantum states, it follows from Schoenberg’s theorem [14] that the function $\kappa_{p}$ (37) is a positive semi-definite (PSD) kernel.
Critical feature of the kernel is that it defines a bound of the dissimilarity of sequences with respect of their probabilistic classification maps (12) as it is demonstrated by the following Proposition.
Proposition 2 (Bound on Predictive Classification Divergence).
Let $\kappa_{p}$ be the predictive generative kernel defined by (37). Then for any two sequences $\mathbf{y}^{1},\mathbf{y}^{2}\in L$ and any horizon $k>0$, the variation distance between their predictive classification maps is bounded by the negative logarithm of the kernel:
$\sup_{c\in C}\left|p_{p}(\mathbf{y}^{1},k,c)-p_{p}(\mathbf{y}^{2},k,c)\right|\leq-C_{k}\log\left(\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})\right).$ (39)
@@
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}]P[c|\mathbf{y^{1}z}])-\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{2}]P[c|\mathbf{y^{2}z}])\bigr{|}$ (41)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq c^{*}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}])\bigr{|}$ (42)
@@
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq-C_{k}\log(\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})),$ (45)
where $c^{*}\leq 1$ is a constant dependent on the particular probabilistic classification map $P[c|\mathbf{yz}]$ and $C_{k}=2c^{*}|\Sigma^{k}|$.
Several predictive quantum kernels can be defined using alternative distance or similarity measures between quantum states, such as the Bures metric and quantum fidelity.
The Bures metric between two quantum states is defined as:
@@
Using the Bures metric, we can define a valid predictive kernel via negative exponentiation:
$\kappa_{p}(\mathbf{y},\mathbf{z})=\exp\left(-\mathcal{B}\bigl{(}\phi(\mathbf{y}),\phi(\mathbf{z})\bigr{)}\right),\quad\mathbf{y},\mathbf{z}\in L^{\mathbf{Q}}.$ (46)
Since the Bures metric is conditionally negative definite (CND), the kernel defined in (46) is positive semi-definite (PSD) by Schoenberg’s theorem.
We can also define a predictive kernel directly from fidelity as follows:
$\kappa_{p}(\mathbf{y},\mathbf{z})=F(\rho_{\mathbf{y}},\rho_{\mathbf{z}}).$ (47)
Fidelity is a symmetric, bounded, and positive-definite similarity measure, and thus the function (47) defines a valid PSD kernel.
@@ 4.2 Structural Generative Kernels
$\kappa_{s}(\mathbf{y},\mathbf{z})=\exp\left(-\mathcal{D}\bigl{(}\phi_{s}(\mathbf{y}),\phi_{s}(\mathbf{z})\bigr{)}\right),\mathbf{y},\mathbf{z}\in L^{Q}$ (49)
@@
$\kappa_{s}(\mathbf{y},\mathbf{z})=\exp\left(-\mathcal{B}\bigl{(}\phi_{s}(\mathbf{y}),\phi_{s}(\mathbf{z})\bigr{)}\right),\mathbf{y},\mathbf{z}\in L^{Q}$ (50)
@@ 5.3 Quantum Kernels vs Classical Kernels
where $\left\lVert(\mathbf{y}^{1}-\mathbf{y}^{2})\right\rVert$ is the Euclidean distance between the sequences, and $\sigma$ is a parameter.
@@ 6 Quantum Kernels in Quantum Circuits Computing Model
In the case of more general feature maps producing mixed states we will use definition 3 and equation (37).